The point (-23, 0) lies on the line -2x + 8y = 46. Determine if this point is a solution of the system
-2x + 8y = 46
x – 2y = -9

Answer by mananth(16949)

x – 2y = -9
plug the values (-23,0) in the above equation
-23-2*0=-9
-23=-9
LHS not equal to RHS
hence not a solution of the system